# An LLM-based Readability Measurement for Unit Tests’ Context-aware Inputs

### v1

@@ 3.2 Judging Component
The Judging Component accepts parameters with contexts and a test as the inputs. For each parameter, it judges whether the test has an input passed to it and matches its context. We utilize various tools to judge the context matching, including LLM, arbitrary NLP tool (this paper uses CoreNLP [39]), and regex matching. We first introduce each of them. Then, we describe combining them to complete the judging process.
Tools.
@@
(2) CoreNLP. Traditional NLP tools (e.g., CoreNLP) can verify the consistency of a character string to the context. CoreNLP supports all context categories of String [39], except Cyberspace (see Table I). Its advantage over LLM is that it is financially cost-free.
(3) Regex matching. Regex matching works for String’s Cyberspace and Number’s all categories. For example, we can use "^0[bB][01]+\$" to check whether a number is binary.
Figure 4: An example of context matching prompts for LLM
@@
Primitive-type input judger. We use different tool combinations for context categories (shown in Table I’s last column).
@@
(4) Rule + Regex matching. This combination checks Number’s Other Formats. Before leveraging regex matching, we add a parameter called BearableLength (this paper set it to $9$). For this category, if an input’s length is shorter than BearableLength, we regard it as readable. Otherwise, we use regex matching to check the inputs: the requirement for “LONGNUMBER” is that an underscore must appear after at most every three consecutive digits (e.g., 11_321_22_745); for “FIXEDLENGTH”, the requirement is that an underscore must exist and appear after a fixed number of consecutive digits (e.g., $1$_1111_1111_1111); for “SCIENTIFIC”, we checks whether the input contains an e or E. We assume the test code is grammatically correct, so as long as a number contains such symbols, it satisfies “SCIENTIFIC”.
@@ 4 EvoSuite Enhanced by C3 (EvoSuiteC3)
The readability contexts mined by the Mining Component of C3 (Sec. 3.1) can not only used by C3’s Judging Component (Sec. 3.2) to judge tests but also can help automated tools generate tests better in readability. In this paper, we propose EvoSuiteC3, i.e., EvoSuite [2] enhanced by C3. Fig. 5 shows its overview. EvoSuite is a search-based test generation tool. It formulates coverage metrics (e.g., Branch Coverage and Weak Mutation [54, 55]) as fitness functions and uses them to guide genetic algorithms [25, 34, 5] to evolve tests until all fitness functions reach optimum or the search budget is out. To enhance EvoSuite, we take different strategies for string and number contexts: For each string context, we provide a group of string seeds (e.g., "Beijing" and "London" for “CITY”) for EvoSuite. These seeds can be randomly selected by EvoSuite to be directly inserted into tests. Secondly, we provide extra fitness functions to guide EvoSuite’s genetic algorithm so that it can generate tests with string values satisfying contexts. Unlike string contexts, number contexts focus on the numbers’ formats instead of values. Hence, we modify the formats of numbers according to contexts after generating tests without any operations in the test generation process.
@@
$\left\{\begin{array}[]{lcl}0&&{isSatisfy(value,}\\ &&{context)=1,}\\ minDistance(value,context)&&{\textnormal{otherwise,}}\\ \end{array}\right.$
@@
$\max{\{f_{c3}(v,c)|(v,c)\in getVCPairs(invocation)\}},$
where $getVCPairs$ returns an invocation’s all value-context pairs. This function returns 00 only if all contexts are satisfied.
@@ 5 Evaluation
The evaluation mainly assesses C3’s performance, i.e., false positive and negative rates, in mining readability contexts from real projects. Additionally, we aim to explore the differences between manual and automated tests under C3 measurement. Moreover, we investigate whether programmers agree with C3. The research questions are as follows:
@@ 5.1.1 Methodology
(1) Subjects. Table II presents the real projects used in this RQ. These projects come from two sources: (1) We select $254$ classes from $100$ projects within the DynaMOSA benchmark [5, 57]; (2) To enrich the project diversity and obtain more classes with manual tests for subsequent RQs, we add $155$ classes from seven open-source projects, namely, Airsonic [58], BroadleafCommerce [59], GoogleMapService [60], OpenMRS [61], OpenRefine [62], PetClinic [63], and Shopizer [64]. In summary, these classes contain $4485$ primitive-type parameters.
5.1.2 Results
@@ 5.1.3 Discussion
(1) LLM’s randomness. We find $154$ ($87\%$) FNs and $49$ ($31\%$) FPs whose projects contain many parameters with similar names, comments, and meanings that are accurately annotated with context, except for them. We assume this is due to LLM’s randomness, which could be mitigated by calling LLM multiple times. We randomly select $30$ cases as a sample. After calling LLM three times, $20$’s readability contexts can be correctly extracted at least one time. After calling LLM five times, the correct cases increase to $24$.
(2) LLM’s insufficient understanding. We find that $28$ ($13\%$) FNs and $102$ ($64\%$) FPs are due to LLM’s insufficient understanding. The first example is a class Playlist from Airsonic [58], which stores music playlists. Its one parameter, name (the playlist’s name), is wrongly recognized as “PERSON”. The second instance is a class Billing from Shopizer [64]. Its parameter, state that could be tagged as “COUNTRY”, is missed by GPT-4-Turbo. We assume that these cases are due to LLM’s insufficient understanding, which could be solved by the LLM with better understanding in the future. On the other hand, our mine-context task can be used as a benchmark to compare different LLMs.
Another issue undermining C3 is that our contexts are incomplete. One example is a class Location from OpenMRS [61]. Its one string parameter, postalCode is not recognized since we do not define “POSTCODE” context (see Sec. 3.1). We do not count these instances as FN, as they fall outside C3’s contexts and exceed its capabilities. To cope with this issue, we allow C3’s users to integrate new contexts into C3 by providing their names, descriptions, and some examples for LLMs (see Sec. 3.1).
@@ 5.2.1 Methodology
(1) Subjects. Table V shows this RQ’s subjects, i.e., $30$ classes with manual tests whose false positive cases have been identified in RQ1 (see Sec. 5.1). We chose these classes because they have manual tests, and we aim to compare manual and automated tests in this RQ.
@@
(2) Baselines. We have five baselines: manual tests, EvoSuite [2], Randoop [3], ChatUniTest [1], and EvoSuiteC3 (Sec. 4). We use EvoSuite and Randoop to represent two traditional test generation techniques, i.e., search-based software testing [25, 2, 34, 5, 35] and random software testing [3, 28, 29, 30]. We use ChatUniTest to represent LLM-based test generation [43, 1, 45] since it is ready-to-use and can continuously guide LLM to generate syntactically correct tests. EvoSuiteC3 is EvoSuite enhanced by C3 (see Sec. 4).
@@
(3) Test generation procedure. For EvoSuite, Randoop, and EvoSuiteC3, we use their default settings to run on each class for producing tests. Due to their randomness, we repeat the procedure for $30$ rounds, and each round’s time budget is two minutes. Besides, we pass all contexts, including false positive ones mined by C3 (Table V), to EvoSuiteC3. For ChatUniTest, we adopt the authors’ settings [1]. For example, the test number for each method under test (MUT) is six, and the LLM is GPT-3.5-Turbo [65]. The only difference is that we use at most ten rounds of calling GPT-3.5-Turbo to generate and repair a test instead of six rounds to obtain more syntactically correct tests. Although GPT-3.5-Turbo also has randomness, we only execute the routine once due to financial costs.
@@
(4) C3 measurement procedure. After the test generation procedure, we employ C3’s judging component (see Sec. 3.2) to evaluate all tests. We execute C3 with GPT-4-Turbo on both manual and automated tests produced by ChatUniTest. Likewise, we utilize GPT-4-Turbo for the first round of tests generated by EvoSuite, Randoop, and EvoSuiteC3. However, considering the substantial model financial costs brought by the tests of 30 rounds, we use GPT-3.5-Turbo to run with C3 for them.
@@
(5) Data analysis. We calculate two metrics for both manual and automated tests generated by ChatUniTest, as well as the first round of EvoSuite and Randoop tests: the number of covered parameters (i.e., the existence of a test that calls the MUT with an input) and parameters with readable inputs (i.e., the existence of a test that calls the MUT with a readable input). Additionally, we show the metric data by true positive contexts and all contexts (see Table V), respectively, to show the effects of false positive cases from C3 on measuring test generation tools. Note that we only show string-type data for all contexts since there are no false positive cases in number-type contexts (see Table V). Furthermore, we compute the average values of these two metrics across all rounds of tests generated by EvoSuite, Randoop, and EvoSuiteC3 to check whether the randomness affects their performance in generating readable inputs.
@@ 5.2.2 Results
(1) String coverage and readability. Table VI presents the performance of five baselines in covered parameters and parameters with readable inputs on $91$ true positive string-type contexts. Overall, four automated tools surpass manual tests in parameter coverage ($86\%$ to $98\%$ versus $48\%$). The readability rates (i.e., the ratio of parameters with readable input to covered parameters) indicate that EvoSuiteC3, ChatUniTest, and manual tests significantly outperform EvoSuite and Randoop in generating readable inputs ($90\%$, $83\%$, and $68\%$ versus $8\%$ and $8\%$). Fig. 1 summarizes test inputs of five baselines.
(2) Number coverage and readability. Table VI also presents baselines’ performance on $24$ number-type contexts. Similarly, four test generation tools (ChatUniTest, EvoSuite, EvoSuiteC3, and Randoop) surpass manual tests in parameter coverage ($88\%$, $71\%$, $71\%$, and $62\%$ versus $50\%$). Their readability rates are close: $92\%$ (manual tests) and $100\%$ (four tools).
(3) Effects of false positive contexts. Table VII shows the results of all string-type contexts (including false positive cases). The parameter coverage delta between all and true positive contexts only exists in Randoop ($86\%$ and $87\%$). The readability deltas range from $1\%$ (EvoSuite and EvoSuiteC3) to $8\%$ (ChatUniTest).
(4) Thirty rounds. We run EvoSuite, Randoop, and EvoSuiteC3 $30$ times to eliminate the randomness. Table VIII shows their average performance. In total, the parameter coverage deltas between one and $30$ rounds range from $0\%$ (EvoSuite’s string) to $4\%$ (EvoSuite’s number). The readability deltas range from $0\%$ (EvoSuiteC3’s number) to $7\%$ (Randoop’s string).
@@ 5.2.3 Discussion
(1) Parameter coverage. The parameter coverage figures show that all test generation tools can generate higher code coverage than manual tests, confirming many studies’ claims and findings [54, 6].
(2) String readability. EvoSuiteC3 and ChatUniTest’s tests surpass manual ones regarding string-input readability, while EvoSuite and Randoop almost fail to generate readable string-type inputs.
@@
(4) Effects of false positive contexts. False positive contexts lead to underestimating the string-type readability performance of ChatUniTest and manual tests. The primary reason is that these two baselines perform well in string-type readability, thus making them sensitive to incorrect contexts. However, such contexts have tiny impact on EvoSuite and Randoop, as these tools fail to generate readable string inputs. Therefore, C3, even with false positive contexts, can be utilized to evaluate traditional tools. Besides, since we pass all contexts to EvoSuiteC3, its performance is nearly identical with or without false positive contexts.
(5) Thirty rounds. The thirty rounds of data confirm that EvoSuite and Randoop stably lag far behind manual tests, ChatUniTest, and EvoSuiteC3 in generating readable string inputs.
Answer to RQ2 ChatUniTest and EvoSuiteC3’s tests surpass manual ones regarding parameter coverage and C3 measurement, while EvoSuite and Randoop almost fail to generate readable string-type inputs.
5.3 RQ3: Programmers’ Readability Measurements

### v2

@@ 3.2 Judging Component
The Judging Component aims to give a judgment for a test’s input readability. It accepts parameters with contexts and a test as the inputs. For each parameter, it judges whether the test has an input passed to it and matches its context. We utilize various tools to judge the context matching as follows. First, we can organize a prompt for an LLM to judge inputs. Second, many NLP tools (e.g., CoreNLP [39] used in this paper) support the Named Entity Recognition (NER) task, i.e., extracting named entities from a piece of unstructured text. Since we use name entities as String’s contexts, we further leverage CoreNLP’s NER functionality to judge string inputs. Third, regex matching can judge inputs if one context has a corresponding regular expression (e.g., “EMAIL”). We first introduce each of them. Then, we describe combining them to complete the judging process.
Figure 4: An example of context matching prompts for LLM
Tools. We introduce the tools to perform context matching as follows.
@@
(2) CoreNLP. We use CoreNLP (a traditional NLP tool) to verify the consistency of a character string to the context via its NER functionality. CoreNLP supports all context categories of String [39], except Cyberspace (see Table I). Its advantage over LLM is that it is financially cost-free.
(3) Regex matching. Regex matching works for String’s Cyberspace and Number’s all categories. For example, we use "^0[bB][01]+\$" to check whether a number is binary.
@@
Primitive-type input judger. We use different tool combinations for context categories (shown in Table I’s last column). We prefer to use multiple tools in combination instead of a single tool to judge inputs because we want to reduce the misjudgment of readable inputs.
@@
(4) Rule + Regex matching. This combination checks the Number’s Other Formats. Before leveraging regex matching, we add a parameter called BearableLength. For this category, if an input’s length is shorter than BearableLength, we regard it as readable. This paper sets BearableLength to $9$, and we will discuss it in Sec. 7. Otherwise, we use regex matching to check the inputs: the requirement for “LONGNUMBER” is that an underscore must appear after at most every three consecutive digits (e.g., 11_321_22_745); for “FIXEDLENGTH”, the requirement is that an underscore must exist and appear after a fixed number of consecutive digits (e.g., $1$_1111_1111_1111); for “SCIENTIFIC”, we checks whether the input contains an e or E. We assume the test code is grammatically correct, so as long as a number contains such symbols, it satisfies “SCIENTIFIC”.
@@ 4 EvoSuite Enhanced by C3 (EvoSuiteC3)
The readability contexts mined by C3’s Mining Component (Sec. 3.1) are not only used by C3’s Judging Component (Sec. 3.2) to judge tests but also help automated tools generate tests with better input readability. We aim to help automated tools generate tests with more readable inputs and consistency in other perspectives (such as coverage) with the original tools. In this paper, we take EvoSuite as an example and propose EvoSuiteC3, i.e., EvoSuite [2] enhanced by C3. We choose EvoSuite instead of Randoop because EvoSuite is a search-based test generation tool. It formulates coverage metrics (e.g., Branch Coverage and Weak Mutation [57, 58]) as fitness functions. It uses them to guide genetic algorithms [25, 34, 5] to evolve tests until all fitness functions reach optimum or the search budget is out. Hence, we integrate C3 into EvoSuite by formulating C3’s contexts as the genetic algorithms’ fitness functions. However, Randoop generates a test by repeating randomly calling the methods of the JAVA class under test until some predefined contract is violated [3], which means that one test generated by Randoop aims to capture one behavior (i.e., a violated contract) of the class under test. Hence, it is difficult to add additional testing objectives to Randoop.
@@
Fig. 5 shows EvoSuiteC3’s overview. Generally, we formulate C3’s string-type contexts as fitness functions to guide EvoSuite’s genetic algorithms to generate readable string-type inputs during the test generation process and modify the formats of numbers to satisfy their parameters’ contexts after the test generation process. Specifically, for each string context, we provide a group of string seeds (e.g., "Beijing" and "London" for “CITY”) for EvoSuite. These seeds are randomly selected by EvoSuite to be directly inserted into tests. Secondly, we provide extra fitness functions to guide EvoSuite’s genetic algorithm in generating tests with string values that satisfy contexts. Unlike string contexts, number contexts focus on the numbers’ formats instead of values. Hence, we modify the formats of numbers according to contexts after generating tests without any operations in the test generation process.
@@
$\left\{\begin{array}[]{lcl}0&&{isSatisfy(value,}\\ &&{context)=1,}\\ minDistance(value,context)&&{\textnormal{otherwise,}}\\ \end{array}\right.$ (1)
@@
$\max{\{f_{c3}(v,c)|(v,c)\in getVCPairs(invocation)\}},$ (2)
where $getVCPairs$ returns an invocation’s all value-context pairs. This function returns the max return values of all the target method’s $f_{c3}$ fitness functions. Hence, it returns 00 if only all $f_{c3}$ fitness functions return 00, i.e., all contexts are satisfied.
@@ 5 Evaluation
The evaluation mainly assesses C3’s performance, i.e., false positive and negative rates, in mining readability contexts from real projects. Additionally, we aim to explore the differences between manual and automated tests under C3 measurement and the improvement brought by EvoSuiteC3 to EvoSuite. Moreover, we investigate whether programmers agree with C3. The research questions are as follows:
[…]
@@ 5.1.1 Methodology
(1) Subjects. Table II presents the real projects used in this RQ. These projects come from two sources: (1) We randomly select $254$ classes from $100$ projects within the DynaMOSA benchmark [5, 60]; (2) To enrich the project diversity and obtain classes with manual tests for subsequent RQs, we add $155$ classes from seven open-source projects, namely, Airsonic [61], BroadleafCommerce [62], GoogleMapService [63], OpenMRS [64], OpenRefine [65], PetClinic [66], and Shopizer [67]. In summary, these classes contain $4485$ primitive-type parameters.
5.1.2 Results
@@ 5.1.3 Discussion
(1) LLM’s randomness. We find $154$ ($87\%$) FNs and $49$ ($31\%$) FPs whose projects contain many parameters with similar names, comments, and meanings that are accurately annotated with context, except for them. We assume this is due to LLM’s randomness, which could be mitigated by calling LLM multiple times. We randomly select $30$ false cases as a sample. After calling LLM three times, $20$’s readability contexts are correctly extracted at least one time. After calling LLM five times, the correct cases increase to $24$.
(2) LLM’s insufficient understanding. We find that $28$ ($13\%$) FNs and $102$ ($64\%$) FPs are due to LLM’s insufficient understanding. One example is a class Playlist from Airsonic [61]. Its one parameter, name (the music playlist’s name), is wrongly recognized as “PERSON”. We assume that these cases are due to LLM’s insufficient understanding, which could be solved by the LLM with better understanding in the future. On the other hand, our mine-context task can be used as a benchmark to compare different LLMs.
Another issue undermining C3 is that our contexts are incomplete. One example is a class Role from OpenMRS [64] (a medical record system). It has one string parameter, role, whose expected readable values include “Medical Student”, “Data Manager”, “Data Assistant”, etc. But it is not recognized since we do not define “Role” context. We do not count these instances as FN, as they fall outside C3’s contexts and exceed its capabilities. To cope with this issue, we allow C3’s users to integrate new contexts into C3. Sec. 6 shows how to add contexts.
@@ 5.2.1 Methodology
(1) Subjects. Table V shows this RQ’s subjects, i.e., $30$ classes with manual tests whose false positive cases have been identified in RQ1 (see Sec. 5.1). These $30$ classes contain $128$ parameters with context detected by C3, of which $115$ were manually confirmed as true positive cases. We chose these classes because they have manual tests, and we aim to compare manual and automated tests in this RQ.
@@
(2) Baselines. We have four baselines: manual tests, EvoSuite [2], Randoop [3], ChatUniTest [1]. We use EvoSuite and Randoop to represent two traditional test generation techniques, i.e., search-based software testing [25, 2, 34, 5, 35] and random software testing [3, 28, 29, 30]. We use ChatUniTest to represent LLM-based test generation [44, 1, 46] since it is ready-to-use and designed to continuously guide LLM in generating syntactically correct tests without manual intervention.
@@
(3) Test generation procedure. For EvoSuite and Randoop, we use their default settings to run on each class for producing tests. Due to their randomness, we repeat the procedure for $30$ rounds, and each round’s time budget is two minutes. For ChatUniTest, we adopt the authors’ settings [1]. For example, the test number for each method under test (MUT) is six, and the LLM is GPT-3.5-Turbo [68]. The only difference is that we use at most ten rounds of calling GPT-3.5-Turbo to generate and repair a test instead of six rounds to obtain more syntactically correct tests. Although GPT-3.5-Turbo also has randomness, we only execute the routine once due to financial costs.
@@
(4) C3 measurement procedure. After the test generation procedure, we employ C3’s judging component (see Sec. 3.2) to evaluate all tests. We execute C3 with GPT-4-Turbo on both manual and automated tests produced by ChatUniTest. Likewise, we utilize GPT-4-Turbo for the first round of tests generated by EvoSuite and Randoop. However, considering the substantial model financial costs brought by running GPT-4-Turbo for the tests of $30$ rounds, we use GPT-3.5-Turbo to run with C3 for them since GPT-3.5-Turbo is much cheaper than GPT-4-Turbo.
@@
(5) Data analysis. We calculate two metrics for both manual and automated tests generated by ChatUniTest, as well as the first round of EvoSuite and Randoop tests: the number of covered parameters (i.e., the existence of a test that calls the MUT with an input) and parameters with readable inputs (i.e., the existence of a test that calls the MUT with a readable input). Additionally, we show the metric data by true positive contexts and all contexts (see Table V), respectively, to show the effects of false positive cases from C3 on measuring test generation tools. Note that we only show string-type data for all contexts since there are no false positive cases in number-type contexts (see Table V). Furthermore, we compute the average values of these two metrics across all rounds of tests generated by EvoSuite and Randoop to check whether the randomness affects their performance in generating readable inputs.
@@ 5.2.2 Results
(1) String coverage and readability. Table VI presents the performance of four baselines (and EvoSuiteC3, see Sec. 5.3) in covered parameters and parameters with readable inputs on $91$ true positive string-type contexts. Overall, three automated tools surpass manual tests in parameter coverage ($86\%$ to $98\%$ versus $48\%$). The readability rates (i.e., the ratio of parameters with readable input to covered parameters) indicate that ChatUniTest and manual tests significantly outperform EvoSuite and Randoop in generating readable inputs ($83\%$ and $68\%$ versus $8\%$ and $8\%$). Fig. 1 summarizes test inputs of four baselines as well as EvoSuiteC3.
(2) Number coverage and readability. Table VI also presents baselines’ performance on $24$ number-type contexts. Similarly, three test generation tools (ChatUniTest, EvoSuite, and Randoop) surpass manual tests in parameter coverage ($88\%$, $71\%$, and $62\%$ versus $50\%$). Their readability rates are close: $92\%$ (manual tests) and $100\%$ (three tools).
(3) Effects of false positive contexts. Table VII shows the results of all string-type contexts (including false positive cases). The parameter coverage delta between all and true positive contexts only exists in Randoop ($86\%$ and $87\%$). The readability deltas range from $1\%$ (EvoSuite) to $8\%$ (ChatUniTest).
(4) Thirty rounds. We run EvoSuite and Randoop $30$ times to eliminate the randomness. Table VIII shows their average performance. In total, the parameter coverage deltas between one and $30$ rounds range from $0\%$ (EvoSuite’s string) to $4\%$ (EvoSuite’s number). The readability deltas range from $1\%$ (EvoSuite’s number) to $7\%$ (Randoop’s string).
@@ 5.2.3 Discussion
(1) Parameter coverage. The parameter coverage figures show that all test generation tools generate higher code coverage than manual tests, confirming many studies’ claims and findings [57, 6].
(2) String readability. ChatUniTest’s tests surpass manual ones regarding string-input readability, while EvoSuite and Randoop almost fail to generate readable string-type inputs.
@@
(4) Effects of false positive contexts. False positive contexts lead to underestimating the string-type readability performance of ChatUniTest and manual tests. The primary reason is that these two baselines perform well in string-type readability, thus making them sensitive to incorrect contexts. However, such contexts have a tiny impact on EvoSuite and Randoop, as these tools fail to generate readable string inputs. Therefore, C3, even with false positive contexts, is sufficient to evaluate traditional tools.
(5) Thirty rounds. The thirty rounds of data confirm that EvoSuite and Randoop lag far behind manual tests and ChatUniTest in generating readable string inputs.
Answer to RQ2 ChatUniTest’s tests surpass manual ones regarding parameter coverage and C3 measurement, while EvoSuite and Randoop almost fail to generate readable string-type inputs.
5.3 RQ3: EvoSuiteC3’s Improvement
Sec. 4 introduces EvoSuiteC3, i.e., EvoSuite enhanced by C3. In this RQ, we aim to determine how much improvement EvoSuiteC3 brings to EvoSuite.
5.3.1 Methodology
(1) Subjects. We use the same subjects as RQ2 (shown in Table V).
(2) Baselines. We mainly use EvoSuite to compare with EvoSuiteC3 to evaluate the improvements brought by EvoSuiteC3. We also use ChatuniTest and manual tests as baselines to compare EvoSuiteC3 since C3 confirmed their excellent input readability in RQ2 (see Sec. 5.2).
(3) Test generation procedure. EvoSuiteC3 is based on EvoSuite and does not introduce any extra running parameters. Hence, we use EvoSuite’s default settings to run EvoSuiteC3 on each class for producing tests. Due to its randomness, we repeat the procedure for $30$ rounds to generate $30$ test suites, and each round’s time budget is two minutes. Besides, we pass all contexts, including false positive ones mined by C3 (Table V), to EvoSuiteC3.
(4) C3 measurement procedure. After the test generation, we use C3’s judging component (see Sec. 3.2) to evaluate EvoSuiteC3’s tests. Like in RQ2, we run C3 with GPT-4-Turbo for the first round of tests generated by EvoSuiteC3. Meanwhile, we use GPT-3.5-Turbo for all rounds of tests since using GPT-4-Turbo for all rounds of tests would cost significant financial costs, and GPT-3.5-Turbo is much cheaper than GPT-4-Turbo.
[…]
5.3.2 Results
(1) Comparison with EvoSuite. The last column of Tables VIVII, and VIII shows the C3 measurement data of EvoSuiteC3. They uniformly show that with other metrics almost the same, the string readability increases from about $10\%$ to around $90\%$. Code 1 presents an example. EvoSuite generates two random inputs for two parameters (username and email), while EvoSuiteC3 generates a person’s name and an email address for them. Fig. 1 confirms that EvoSuiteC3 generates readable inputs while EvoSuite’s inputs are commonly random.
(2) Comparison with Manual Tests and ChatUniTest. Table VI presents the C3 measurement data of manual tests, ChatUniTest, and EvoSuiteC3. First, ChatUniTest and EvoSuiteC3 surpass manual tests in both parameter coverage and string readability. Second, EvoSuiteC3 reaches the highest string readability for all classes ($90\%$). EvoSuiteC3 only performs worse than manual tests in the GoogleMapService project’s string readability score ($76\%$ versus $83\%$).
[…]
Mean Median Mean Median
1 30 93.5% 100% 89.1% 100%
2 7 94% 100% 78% 75%
3 3 99.8% 100% 73.7% 75%
[…]
5.3.3 Discussion
[…]
Answer to RQ3 Under C3 measurement, EvoSuiteC3 significantly improves EvoSuite’s string input readability by satisfying from about $10\%$ of contexts to around $90\%$.
5.4 RQ4: Programmers’ Readability Measurements
In this RQ, we aim to determine whether programmers agree with C3’s measurement of tests’ readability.
5.4.1 Methodology